# TimeScope: Towards Task-Oriented Temporal Grounding In Long Videos

### v2

@@ B.1 IoU and mIoU
IoU@0.3: Percentage of predictions with IoU 0.3
IoU@0.5: Percentage of predictions with IoU 0.5
IoU@0.7: Percentage of predictions with IoU 0.7
B.2 Query Center Robustness
@@
Appendix C Model Arch and Training detail
Timescope initializes with the base model of VideoXL2 because it can handle very long sequences, facilitating the construction of long video Task-oriented Temporal Grounding tasks. The overall architecture is shown in the figure above. The architecture of Timescope consists of four components: a visual encoder, Dynamic Token Synthesis (DTS), an MLP projector, and a Large Language Model (LLM). For the visual encoder, we use SigLIP to encode visual inputs into dense visual features. The DTS module, located after the visual encoder, is constructed by combining spatial-temporal attention blocks and 3D convolutional layers. It aims to process visual features extracted from four consecutive visual inputs as a single group. This design has been validated in Video-XL-Pro. Next, the architecture applies average pooling to adjacent features to further compress the representation. Then, a two-layer MLP projector processes these pooled features, projecting them into the embedding space of the LLM. For our LLM, we adopt Qwen2.5-7B.
Furthermore, in Timescope, to enhance the model’s temporal understanding capability, we prepend explicit timestamp tokens (e.g., Time: 4.0 Second) for every four consecutive frames group across the entire frames sequence. This direct timing information substantially improves the model’s temporal awareness. During training, we randomly perform shift operations (e.g., 4s to 1004 seconds) and cut operations (e.g., cutting out 10-20 seconds of a video as a single data point) on the explicit timestamp tokens of the input video as a whole. This training technique greatly enhances Timescope’s multi-stage understanding capability and robustness to query positions, as shown in Table B.2.
Appendix D Experimental Settings & Additional Results
We elaborate on the training and inference details of Timescope. The reported hyperparameters cover Stage 1, and 2, as specified in Table 1. For the inference details, we emphasize the particular context length for different benchmarks, as shown in Table 2.
@@ Appendix D Experimental Settings & Additional Results
Appendix E VideoQA Task Details
E.1 VideoQA with Temporal Grounding
@@
E.2 Prompt template for VideoQA
@@
Appendix F Limitations & Future Work
@@
Appendix G Broader Impacts Statement
@@
Appendix H Qualitative Results
In Figure 2–4, we present the qualitative results of Timescope on ToTG-bench and V-STaR. Timescope demonstrates accurate understanding of questions and the ability to provide temporal grounding. In V-STaR, we show the results when two-stage reasoning is applied, and it can be seen that Timescope exhibits robust performance with good coarse-grained segment retrieval and fine-grained temporal grounding capabilities.

### v3

@@ D.1 IoU and mIoU
IoU@0.3: Percentage of predictions with IoU 0.3
IoU@0.5: Percentage of predictions with IoU 0.5
IoU@0.7: Percentage of predictions with IoU 0.7
D.2 Query Center Robustness
Appendix E Experimental Settings & Additional Results
Appendix F VideoQA Task Details
F.1 VideoQA with Temporal Grounding
@@
F.2 Prompt template for VideoQA
@@
Appendix G Limitations & Future Work
@@
Appendix H Broader Impacts Statement
@@
Appendix I Qualitative Results
In Figure 5–7, we present the qualitative results of Timescope on ToTG-bench and V-STaR. Timescope demonstrates accurate understanding of questions and the ability to provide temporal grounding. In V-STaR, we show the results when two-stage reasoning is applied, and it can be seen that Timescope exhibits robust performance with good coarse-grained segment retrieval and fine-grained temporal grounding capabilities.